Jefferson Lab HALL A Logbook for February 05 - All EXCEPT Automatic Entries

  • Main INDEX

    Sunday

    
    

    Friday

  • 140736 02/25/05 14:24 saw Scaler clear for floor configuration
  • Thursday

  • 140701 02/24/05 13:06 atrig Trigger-Download
  • Wednesday

    
    

    Tuesday

    
    

    Thursday

  • 140636 02/17/05 16:15 Olivier Swapped delay cables for EL11-12 ADC
  • Wednesday

  • 140633 02/09/05 15:03 saw New scaler channels and new scaler display
  • Tuesday

    
    

    Monday

    
    

    Thursday

  • 140626 02/03/05 10:13 Olivier Window widths for T6 trigger
  • Wednesday

  • 140620 02/02/05 12:09 Olivier Back to SRC configuration after N20
  • 140619 02/02/05 11:40 Olivier T6 trigger formation and overlap
  • 140618 02/02/05 11:05 Kolarkar handing Hall A controls to Mike Seely
  • 140614 02/02/05 09:52 Kolarkar opened JT valve
  • 140613 02/02/05 09:40 Holmstrom N20 Over
  • 140603 02/02/05 08:16 E.Chudakov Run with BigBite
  • 140601 02/02/05 08:11 E.Chudakov Shift summary - owl
  • 140592 02/02/05 05:40 E.Chudakov HRSL settings
  • 140590 02/02/05 04:58 Kelleher target to LH2
  • 140589 02/02/05 04:56 E.Chudakov HRSL setting
  • 140584 02/02/05 03:12 Kelleher modified N20 run plan
  • 140583 02/02/05 03:09 Kelleher target to empty
  • 140582 02/02/05 02:51 E.Chudakov HRSL polarity change
  • 140579 02/02/05 01:23 Kelleher Target to 4cm LD2 position
  • 140577 02/02/05 00:25 Kelleher N20 Run Plan
  • Tuesday

  • 140573 02/01/05 23:58 Annand Shift Summary
  • 140572 02/01/05 23:54 Kolarkar N20 veto settings/analysis (more readable)
  • 140567 02/01/05 22:16 Kolarkar N20 veto settings/analysis
  • 140559 02/01/05 18:35 Kolarkar BB HV turned off
  • 140556 02/01/05 18:25 Kolarkar rates screen snapshot for run # 2276
  • 140552 02/01/05 18:08 Kolarkar BB rates screen
  • 140549 02/01/05 15:51 Olivier Thresholds on BB
  • 140537 02/01/05 11:11 R. Michaels DAQ for n20 test
  • 140534 02/01/05 10:56 L. Zhu target warm up
  • 140533 02/01/05 10:36 Olivier Top-right, not bottom-right delay box
  • 140532 02/01/05 10:30 Olivier Current T7 delay scheme for SRC running
  • 140531 02/01/05 10:29 R. Feuerbach ESR Helium flow dropped
  • 140530 02/01/05 10:25 Olivier( ND scaler setup for Lingyan's BB rate test
  • 140527 02/01/05 09:02 P. Monaghan Owl Shift Summary - until 9am
  • 140512 02/01/05 06:46 ran HV lists for BB and N-array
  • 140507 02/01/05 06:00 P. Monaghan MCC RF cavity problem
  • 140504 02/01/05 05:48 ran ADC of N array in(top) and out(buttom) of the TOF peak
  • 140493 02/01/05 03:38 ran moder art ..
  • 140492 02/01/05 03:29 ran modern art
  • 140479 02/01/05 01:47 ran TOF to N-array,with no adc cut